# Deep Linear Hawkes Processes

### v1

@@ 3.1 Continuous-Time Latent Linear Hawkes Layer
We first start by generalizing the intensity of the linear Hawkes process, Eq. 6:
$\mathrm{d}\bm{\lambda}_{t}$ $=-\bm{\beta}(\bm{\lambda}_{t-}-\bm{\nu}_{t})\mathrm{d}t+\bm{\alpha}\mathrm{d}\mathbf{N}_{t}=-\bm{\beta}\bm{\lambda}_{t-}\mathrm{d}t+\bm{\beta}\bm{\nu}_{t}\mathrm{d}t+\bm{\alpha}\mathrm{d}\mathbf{N}_{t},$ (9)
whereby the background intensity $\bm{\nu}_{t}$ is allowed to vary over time. If we compare this to the recurrence in Eq. 7, we see that the intensity in the LHP, $\bm{\lambda}_{t}$ controlled by decay rates $\bm{\beta}$, is analogous the state in the linear SSM, $\mathbf{x}(t)$ controlled by state matrix $\mathbf{A}$. Additionally, the time-varying baseline intensity in the LHP, $\bm{\nu}_{t}$, is analogous to the SSM input signal, $\mathbf{u}(t)$. What is unique to the LHP is the (mark-specific) impulse $\bm{\alpha}\mathrm{d}\mathbf{N}_{t}$. This impulse is important because it allows the model to instantaneously incorporate information from events as they occur, introducing discontinuities in the output signals of the otherwise continuously-integrated system, unlike conventional SSMs.
With this in mind, we adapt Eq. 9 such that it can replace the typical state-space recurrence in an SSM, Eq. 7. To do so, we replace the non-negative $\bm{\beta}$ with an unrestricted state matrix $\mathbf{A}\in\mathbb{R}^{P\times P}$. Next, given an input signal $\mathbf{u}_{t}\in\mathbb{R}^{H}$ we project it to $P$ dimensions with an input matrix $\mathbf{B}\in\mathbb{R}^{P\times H}$ to replace $\bm{\nu}_{t}$. [^2] Here, we index time $t$ via subscripts (e.g., $\mathbf{u}_{t}$) rather than an argument $\left(\mathbf{u}(t)\right)$ to emphasize that these are stochastic (jump) processes rather than deterministic functions. What was originally the intensity $\bm{\lambda}_{t}$ is now relabeled to be the state of the layer $\mathbf{x}_{t}$. Finally, we allow the impulses to be low-rank by having a shared set of mark embeddings $\bm{\alpha}\in\mathbb{R}^{R\times K}$ with rank $R$ that are brought into $P$ dimensions with a layer-specific embedding matrix $\mathbf{E}\in\mathbb{R}^{P\times R}$. For simplicity, we set $R=H$ for all our experiments. The equation for the output signal $\mathbf{y}_{t}$ is left unchanged from Eq. 8, where $\mathbf{C}\in\mathbb{R}^{H\times P}$ and $\mathbf{D}\in\mathbb{R}^{H\times H}$. All of this results in the set of equations that makes up what we call the latent linear Hawkes layer:
$\mathrm{d}\mathbf{x}_{t}$ $=-\mathbf{Ax}_{t-}\mathrm{d}t+\mathbf{ABu}_{t-}\mathrm{d}t+\mathbf{E}\bm{\alpha}\mathrm{d}\mathbf{N}_{t}$ (10)
$\mathbf{y}_{t}$ $=\mathbf{Cx}_{t}+\mathbf{Du}_{t},$ (11)
where the initial state $\mathbf{x}_{0}\in\mathbb{R}^{P}$ is learned. Realizations of this layer can be seen in Fig. 2.
3.2 Continuous-Time Deep Linear Hawkes Process Architecture
Inspired by deep SSMs, our MTPP is formed by stacking LLH layers, chaining the output signal $\mathbf{y}$ of one layer to the input $\mathbf{u}$ of another with non-linear transforms in between. The final layer’s output is then transformed into the intensity $\bm{\lambda}$. An illustration of the DLHP architecture is shown in Fig. 2.
Let $L$ be the number of desired LLH layers that comprise a DLHP with input and output signals $\mathbf{u}^{(l)}$ and $\mathbf{y}^{(l)}$ respectively for layers $l=1,\dots,L$. For the very first layer, the only input available to condition on are the event occurrences themselves. As such, we set $\mathbf{u}_{t}^{(1)}=\mathbf{0}$ for all $t\geq 0$.
In general, a layer’s output $\mathbf{y}^{(l)}:=\text{LLH}^{(l)}(\mathbf{u}^{(l)},\mathcal{H})$ is passed into a non-linear activation function $f$ (we use $f(z):=\text{GELU}(z)$ (Hendrycks & Gimpel, 2016)), summed with the residual stream $\mathbf{u}^{(l)}$, and normalized with LayerNorm (Ba, 2016) to compute the next layer’s input. More formally,
$\mathbf{u}^{(l+1)}_{t}:=\text{LayerNorm}^{(l)}(f(\mathbf{y}^{(l)}_{t})+\mathbf{u}^{(l)}_{t})$ (12)
for $t\geq 0$ and $l=1,\dots,L$. We use the same strategies for initialization as S5 (Smith et al., 2022), based off the performant HiPPO initialization scheme (Gu et al., 2020). Due to the transformations, unlike the original LHP, we cannot guarantee the output of the final layer is positive. Therefore, similar to Mei & Eisner (2017), we apply an affine projection followed by a rectifying transformation to enforce non-negative intensity:
$\bm{\lambda}_{t}:=\mathbf{s}\odot\text{softplus}((\mathbf{W}\mathbf{u}_{t-}^{(L+1)}+\mathbf{b})\odot\mathbf{s}^{-1})$ (13)
for $t\geq 0$ and where $\mathbf{W}\in\mathbb{R}^{K\times H}$, $\mathbf{b},\log(\mathbf{s})\in\mathbb{R}^{K}$, and $\odot$ is an element-wise product. Eq. 13 implements the “Proj. & Softplus” layer in Fig. 2. The intensity at time $t$ always uses the left-limit of $\mathbf{u}^{(L+1)}$, which in turn uses the left-limit of $\mathbf{y}^{(l)}$ and $\mathbf{u}^{(l)}$ for all $l$ to ensure that it has no information of any events that may or may not have occurred at time $t$ is used.
@@ 3.2 Continuous-Time Deep Linear Hawkes Process Architecture
On the Relationship With the Linear Hawkes Process Before discussing how to compute the DLHP and its variations, we briefly reflect on the relationship between the DLHP and LHP. The derivation presented above shows the steps to modify an LHP to be a deep SSM. The connection, parameterization, and equivalence we explore does not materially affect the implementation; this was intended to concretely define how our model differs from a classical model, and to retain the intuition from the simpler LHP (even if the direct interpretability of the LHP is somewhat lost). Alternatively, one could have simply attempted to convert a deep SSM into an MTPP. Likewise, the steps taken and the result would be similar; however, the relationship to other models would be markedly less clear. We hope our exposition makes it clear how the DLHP is a natural extension of a known and well-used model rather than an arbitrary modification to a recent architecture.
3.3 Discretizing & Diagonalizing the LLH Layer
@@
Unlike the LHP intensity, the recurrence in the LLH layer does not permit an analytic solution. As such, we must discretize the continuous-time process to compute values of the layer at specified time points. If we approximate the input signal by treating it as constant over an update interval, also known as a zero-order hold (ZOH) assumption (Iserles, 2009), then we can achieve a closed-form exact update to the recurrence relation. However, unfortunately, this involves a computationally-expensive matrix exponential in the update rule. To circumvent this, we first diagonalize the system and then impose the zero-order hold restriction on it. Doing so converts the matrix exponential into an element-wise exponential operation. This is done for all LLH layers that compose the DLHP. Note that this is same general approach taken by Smith et al. (2022) for deep SSMs.
@@
Diagonalization Let $-\mathbf{A}$ be diagonalizable with a factorization of $\mathbf{V}\bm{\Lambda}\mathbf{V}^{-1}$, where $\mathbf{V},\bm{\Lambda}\in\mathbb{C}^{P\times P}$ and $\bm{\Lambda}$ is a diagonal matrix of eigenvalues. An equivalent, diagonalized LLH is then
@@
where $\tilde{\mathbf{x}}_{t}=\mathbf{V}^{-1}\mathbf{x}_{t}$, $\tilde{\mathbf{B}}=-\mathbf{V}^{-1}\mathbf{B}$, $\tilde{\mathbf{E}}=\mathbf{V}^{-1}\mathbf{E}$, and $\tilde{\mathbf{C}}=\mathbf{CV}$. Note that in practice we directly parameterize $\tilde{\mathbf{B}}$, $\tilde{\mathbf{C}}$, and $\tilde{\mathbf{E}}$ to avoid having to learn and invert $\mathbf{V}$. The eigenvalues $\bm{\Lambda}$ are also directly parameterized and constrained with negative real-components for stability (Davis, 2013). While the dynamics are diagonalized, we note this does not mean that we are independently modeling the intensities of different mark types. This can be seen two ways: First, the diagonalized dynamics are equivalent to the original dynamics (see Eq. 10, given the system can be diagonalized on the complex plane). Alternatively, the marks interact through the dense input and output matrices, the position-wise non-linearity, the mark embeddings, and the final intensity rectification layer.
Discretization We then employ a ZOH discretization to create a closed-form update from the diagonalized continuous-time system. The ZOH assumption holds the input $\mathbf{u}$ constant over the integration period. This results in the following update rule that transitions from $\mathbf{x}_{t}$ to $\mathbf{x}_{t^{\prime}}$, where, by construction, no events occur in $(t,t^{\prime})$:
$\tilde{\mathbf{x}}_{t^{\prime}}$ :=$:=\begin{cases}\bar{\bm{\Lambda}}\tilde{\mathbf{x}}_{t}+(\bar{\bm{\Lambda}}-\mathbf{I})\tilde{\mathbf{B}}\mathbf{u}_{t^{\prime}-}&\text{ if no event at }t^{\prime}\\ \bar{\bm{\Lambda}}\tilde{\mathbf{x}}_{t}+(\bar{\bm{\Lambda}}-\mathbf{I})\tilde{\mathbf{B}}\mathbf{u}_{t^{\prime}-}+\tilde{\mathbf{E}}\bm{\alpha}_{k}&\text{ if event of type }k\text{ at }t^{\prime}\end{cases}$ (16)
where $\bar{\bm{\Lambda}}:=\exp(\bm{\Lambda}(t^{\prime}-t))$ (derivation in Section B.2). Please refer to Fig. 2 for an illustration.
Note that the ZOH is an exact update when $\mathbf{u}$ is constant over the window $[t,t^{\prime})$. While we choose the constant value to set $\mathbf{u}$ to be as $\mathbf{u}_{t^{\prime}-}$, it is worth noting that technically any value $\mathbf{u}_{s}$ for $s\in[t,t^{\prime})$ is valid. We explore this design decision and the impact it has on performance in more detail in Sections B.4 and D.4. It is important that $\mathbf{u}_{t^{\prime}}$ is not used as the ZOH value to avoid data leakage.
@@ 3.4 Input-Dependent Dynamics
Inspired by recent developments in modern SSMs (e.g., Mamba (Gu & Dao, 2023)), we also consider allowing the dynamics of the system to vary depending on the input and history of previous events. This can allow for more expressive intensities. For instance, dynamically adjusting the real components of $\bm{\Lambda}$ to be smaller will result in more influence from history. Alternatively, larger values will result in more quickly “forgetting” the influence of previous events for a given hidden state channel. This is formalized with the following recurrence relation:
@@
for $t\in(t_{i},t_{i+1}]$ where $\bm{\Lambda}_{i}:=\text{diag}\left(\text{softplus}(\mathbf{W}^{\prime}\mathbf{u}_{t_{i}}+\mathbf{b}^{\prime})\right)\bm{\Lambda}$ with $\mathbf{W}^{\prime}\in\mathbb{R}^{P\times H}$ and $\mathbf{b}^{\prime}\in\mathbb{R}^{P}$. Note that this is still conditionally linear in time as even though $\bm{\Lambda}_{i}$ changes it is entirely input-dependent based on $\mathbf{u}$ and not dependent on previous values of $\mathbf{x}$.
3.5 Computing LLH Recurrence
Thus far, we have created the LLH layer by diagonalizing and discretizing a modified SSM. As discussed earlier, we would like to take advantage of the efficient parallel scans leveraged by many SSM-based models (Smith et al., 2022; Gu & Dao, 2023; Dao & Gu, 2024). Below we explain how we can still use the parallel scan, despite the modified recurrence.
Parallel scans admit efficient inference over linear recurrences of the form $\mathbf{z}_{i+1}=\mathbf{A}_{i}\mathbf{z}_{i}+\mathbf{b}_{i}$ (Blelloch, 1990). Although we have added an impulse to the recurrence, this is still intrinsically of this form, where $\mathbf{z}_{i}:=\mathbf{x}_{t_{i}}$, $\mathbf{A}_{i}:=\exp(\bm{\Lambda}_{i}(t_{i+1}-t_{i}))$, and $\mathbf{b}_{i}:=(\mathbf{A}_{i}-\mathbf{I})\tilde{\mathbf{B}}\mathbf{u}_{t_{i+1}-}+\tilde{\mathbf{E}}\bm{\alpha}_{k_{i+1}}$. As a result, we can leverage efficient parallel scans to compute the sequence of right-limits $\mathbf{x}_{t_{1:N}}$ in parallel across the sequence length. The corresponding left-limits $\mathbf{x}_{t_{1:N}-}$ can then be efficiently computed after by subtracting off $\tilde{\mathbf{E}}\bm{\alpha}_{k_{1:N}}$ from $\mathbf{x}_{t_{1:N}}$. In Algorithms 1, 2 and 3 we compactly detail how to use a parallel scan to compute the sequence of right limits given events; how to evolve those right limits to compute left limits; and then how to subsequently compute the log-likelihood of the sequence.
4 Related Works

### v2

@@ 3.1 Continuous-Time Latent Linear Hawkes Layer
To develop a recurrent layer for an MTPP model, we start by reviewing the equations for an LHP intensity, Eq.˜3, and an SSM state, Eq.˜4. If we allow $\bm{\nu}$ in the LHP to vary over time, we obtain:
@@
[…]
When presented together the parallels between them become apparent: the LHP intensity, $\color[rgb]{1,0,0}\bm{\lambda}_{t}$} controlled by decay rates $\color[rgb]{1,0.49,0}\bm{\beta}$}, is analogous to the state in the linear SSM, $\color[rgb]{1,0,0}\mathbf{x}(t)$} controlled by state matrix $\color[rgb]{1,0.49,0}\mathbf{A}$}. The time-varying background LHP intensity, $\color[rgb]{0.55,0.14,1}\bm{\nu}_{t}$}, is analogous to the SSM input signal after being projected into the state-space, $\color[rgb]{0.55,0.14,1}\mathbf{Bu}(t)$}. Compared to SSMs which allow for recurrence over a latent space $\mathbb{R}^{P}$, Eq.˜6 is limited in expressivity due to restricted dynamics $\color[rgb]{1,0.49,0}\bm{\beta}$} and fixed dimensionality tied to the number of possible marks $K$. Not present in Eq.˜7 is an impulse $\color[rgb]{0.55,0,0}\bm{\alpha}\mathrm{d}\mathbf{N}_{t}$}, which is crucial for allowing the recurrence to condition on abrupt event occurrences over time.
Figure 2: Schematic of our state-space point process. We depict the internals of a single LLH layer of the model in their continuous time form (left) and as discrete computations (right). Black arrows can be concurrently computed in logarithmic time, and gray arrows in constant time.
With this in mind, we combine Eqs.˜6 and 7 to resolve their respective deficiencies. This results in the following set of stochastic jump differential equations that we call the latent linear Hawkes (LLH) layer, which will serve as a continuous-time recurrence layer in our proposed model:
LLH: $\mathrm{d}{\color[rgb]{1,0,0}\mathbf{x}_{t}}$ $=-{\color[rgb]{1,0.49,0}\mathbf{A}}{\color[rgb]{1,0,0}\mathbf{x}_{t-}}\mathrm{d}t+{\color[rgb]{1,0.49,0}\mathbf{A}}{\color[rgb]{0.55,0.14,1}\mathbf{Bu}_{t-}}\mathrm{d}t+{\color[rgb]{0.55,0,0}\mathbf{E}\bm{\alpha}\mathrm{d}\mathbf{N}_{t}}$ $\quad\in\mathbb{R}^{P}$ (8)
$\mathbf{y}_{t}$ $=\mathbf{C}{\color[rgb]{1,0,0}\mathbf{x}_{t}}+\mathbf{D}{\color[rgb]{0.55,0.14,1}\mathbf{u}_{t}}$ $\quad\in\mathbb{R}^{H}$ (9)
where ${\color[rgb]{0.55,0,0}\bm{\alpha}}\in\mathbb{R}^{R\times K}$, ${\color[rgb]{0.55,0,0}\mathbf{E}}\in\mathbb{R}^{P\times R}$, and the standard SSM parameters share the same dimensionality as in Eqs.˜4 and 5. Note that the resulting dynamics within the LLH layer are much more expressive than the LHP that it was inspired by, due to both operating in a separate set of latent dimensions and by having a general dynamics matrix ${\color[rgb]{1,0.49,0}\mathbf{A}}\in\mathbb{R}^{P\times P}$, compared to the more restrictive $\exp(-{\color[rgb]{1,0.49,0}\bm{\beta}})\in\mathbb{R}^{K\times K}_{\geq 0}$. For parameter efficiency, we allow the new impulse term $\color[rgb]{0.55,0,0}\mathbf{E}\bm{\alpha}\mathrm{d}\mathbf{N}_{t}$} to be computed as a product of model-wide, low-rank mark embeddings $\color[rgb]{0.55,0,0}\bm{\alpha}$} of rank $R$ and a layer-specific projection into state-space $\color[rgb]{0.55,0,0}\mathbf{E}$}. For simplicity, we set $R:=H$ in practice. Realizations of this layer are shown in Fig.˜2.
This LLH will serve as the main form of recurrence in a larger continuous-time model that we define in Section˜3.4. In deriving this layer as an extension from LHPs, we retain a strong inductive bias towards event sequences. Furthermore, by drawing connections between this layer and traditional state-space equations, we allow the downstream model to benefit from many of the innovations designed for deep SSMs in recent years, such as parallel computation.
3.2 Diagonalizing, Discretizing & Computing the LLH Recurrence
@@
Unlike the LHP intensity, there is no analytical solution to the continuous-time LLH recurrence due to the continuously-integrated $\mathbf{u}(t)$ signal. We must therefore approximate the system at specific timepoints. If we approximate the input signal by treating it as constant over an update interval, also known as a zero-order hold (ZOH) assumption (Iserles, 2009), then we can achieve a closed-form exact update to the recurrence relation. To avoid a computationally expensive matrix exponential in the update rule, we adopt the same general approach as Smith et al. (2022) for deep SSMs by first diagonalizing the system and then imposing the zero-order hold restriction on it. This converts the matrix exponential into an element-wise exponential operation for each LLH layer in S2P2.
Diagonalization:
@@
Let $\mathbf{A}$ be diagonalizable with a factorization of $\mathbf{V}\bm{\Lambda}\mathbf{V}^{-1}$, where $\mathbf{V},\bm{\Lambda}\in\mathbb{C}^{P\times P}$ and $\bm{\Lambda}$ is a diagonal matrix of eigenvalues. An equivalent, diagonalized LLH is then
@@
where $\tilde{\mathbf{x}}_{t}=\mathbf{V}^{-1}\mathbf{x}_{t}$, $\tilde{\mathbf{B}}=-\mathbf{V}^{-1}\mathbf{B}$, $\tilde{\mathbf{E}}=\mathbf{V}^{-1}\mathbf{E}$, and $\tilde{\mathbf{C}}=\mathbf{CV}$. We can then directly parameterize $\tilde{\mathbf{B}}$, $\tilde{\mathbf{C}}$, and $\tilde{\mathbf{E}}$ to avoid learning and inverting $\mathbf{V}$. We use the same initialization strategies as S5 (Smith et al., 2022) for $\tilde{\mathbf{B}}$, $\tilde{\mathbf{C}}$, and $\tilde{\mathbf{E}}$, based off the HiPPO initialization scheme (Gu et al., 2020). The eigenvalues $\bm{\Lambda}$ are also directly parameterized and constrained with negative real-components to enforce stability (Davis, 2013). While the dynamics are diagonalized, we note this does not mean that we are modeling the intensities of different mark types as independent. Marks interact through both the position-wise nonlinearities and the learned projections $\tilde{\mathbf{B}}$, $\tilde{\mathbf{C}}$ and $\tilde{\mathbf{E}}$ (since the diagonalized dynamics are equivalent to the original dynamics in Eq.˜8, given that the system can be diagonalized on the complex plane).
Discretization:
We then use a ZOH discretization to create a closed-form update from the diagonalized continuous-time system; please refer to Fig.˜2 for an illustration and Section˜B.2 for the full derivation. This results in the following update rule in the diagonalized eigenbasis, that transitions from $\mathbf{x}_{t}$ to $\mathbf{x}_{t^{\prime}}$ without a matrix exponential, where, by construction, no events occur in $(t,t^{\prime})$:
[…]
where $\bar{\bm{\Lambda}}:=\exp(\bm{\Lambda}(t^{\prime}-t))$. ZOH is an exact update when $\mathbf{u}$ is constant over the window $t,t^{\prime})$; as such, we can choose any value $\mathbf{u}_{s}$ for $s\in[t,t^{\prime})$ to hold constant as the input over the integration period. We opt to use $\mathbf{u}_{t^{\prime}-}$ so that the model can condition on the fact that no events have occurred between $t$ and $t^{\prime}$. This design decision and its impact on performance are explored in more detail in Sections˜B.4 and D.6. It is important that $\mathbf{u}_{t^{\prime}}$ is not used as the ZOH value to avoid data leakage.
Computing LLH Recurrence:
The final component is to derive how to use parallel scans to efficiently evaluate the closed-form updates for the modified LLH recurrence in parallel. Parallel scans admit efficient inference over linear recurrences of the form $\mathbf{z}_{i+1}=\mathbf{R}_{i}\mathbf{z}_{i}+\mathbf{b}_{i}$ (Blelloch, 1990). Although we have an impulse in the recurrence, Eq.˜8 is still intrinsically of this form, where $\mathbf{z}_{i}:=\mathbf{x}_{t_{i}}$, $\mathbf{R}_{i}:=\exp(\bm{\Lambda}_{i}(t_{i+1}-t_{i}))$, and $\mathbf{b}_{i}:=(\mathbf{R}_{i}-\mathbf{I})\tilde{\mathbf{B}}\mathbf{u}_{t_{i+1}-}+\tilde{\mathbf{E}}\bm{\alpha}_{k_{i+1}}$. As a result, we can leverage parallel scans to compute the sequence of right-limits $\mathbf{x}_{t_{1:N}}$ in parallel across the sequence length. The corresponding left-limits $\mathbf{x}_{t_{1:N}-}$, which will later be used to calculate event intensities, can then be directly and efficiently computed by subtracting the impulse, $\tilde{\mathbf{E}}\bm{\alpha}_{k_{1:N}}$, from $\mathbf{x}_{t_{1:N}}$.
@@ 3.3 Input-Dependent Dynamics
Inspired by recent developments in modern SSMs (e.g., Mamba (Gu and Dao, 2023)), we also consider allowing the dynamics of the system to vary depending on the input and on the history of previous events. This can allow for more expressive intensities. For instance, dynamically adjusting the real components of $\bm{\Lambda}$ to be smaller will result in more influence from history. Alternatively, larger values will result in more quickly “forgetting” the influence of previous events for a given hidden state channel. This is formalized with the following recurrence relation for $t\in(t_{i},t_{i+1}]$:
$\mathrm{d}\tilde{\mathbf{x}}_{t}$ $:=\bm{\Lambda}_{i}\tilde{\mathbf{x}}_{t-}\mathrm{d}t+\bm{\Lambda}_{i}\tilde{\mathbf{B}}\mathbf{u}_{t-}\mathrm{d}t+\tilde{\mathbf{E}}\bm{\alpha}\mathrm{d}\mathbf{N}_{t},$ (13)
@@
where $\bm{\Lambda}_{i}:=\text{diag}\left(\text{softplus}(\mathbf{W}^{\prime}\mathbf{u}_{t_{i}}+\mathbf{b}^{\prime})\right)\bm{\Lambda}$ with $\mathbf{W}^{\prime}\in\mathbb{R}^{P\times H}$ and $\mathbf{b}^{\prime}\in\mathbb{R}^{P}$. This is conditionally linear in time, as even though $\bm{\Lambda}_{i}$ changes, it is entirely input-dependent based on $\mathbf{u}$ and not dependent on previous values of $\mathbf{x}$, and hence we can still use parallel scans as discussed above.
3.4 State-Space Point Process Architecture
We are now well-positioned to present the state-space point process (S2P2), a flexible and parallel deep continuous-time model for marked intensities $\bm{\lambda}_{t}$. This model is shown in Figs.˜1(a) and 2.
We have demonstrated how to define an efficient, parallelizable and scalable core MTPP layer in the LLH layer. While the diagonalized and discretized LLH layer is more expressive than the LHP, it remains fundamentally linear. To compensate, we take inspiration from deep SSMs and alternate $L$ LLH layers with position-wise nonlinearities. This creates a nonlinear model that can still use parallel computation over the sequence length, but is highly expressive compared to each linear recurrence.
We have an input signal in two parts: (i) the continuously-integrated signal $\mathbf{u}_{t}$ and (ii) the discrete event impulses $\bm{\alpha}_{k}$. For the first layer, the only inputs available to condition on are the event impulses themselves, so we set $\mathbf{u}_{t}^{(1)}=\mathbf{0}$ for all $t\geq 0$. At deeper layers, we have a layer-specific impulse as well as the continuously integrated signal from the previous layer. In general, a layer’s output $\mathbf{y}^{(l)}:=\text{LLH}^{(l)}(\mathbf{u}^{(l)},\mathcal{H})$ is passed into a nonlinear activation function $\sigma$ (we use $\sigma(z):=\text{GELU}(z)$ (Hendrycks and Gimpel, 2016)), summed with the residual stream $\mathbf{u}^{(l)}$, and normalized with LayerNorm (Ba, 2016) for the next layer’s input. Formally, for $t\geq 0$ and $l=1,\dots,L$, then
$\mathbf{u}^{(l+1)}_{t}:=\text{LayerNorm}^{(l)}(\sigma(\mathbf{y}^{(l)}_{t})+\mathbf{u}^{(l)}_{t})$. (14)
Due to the unrestricted nature of the recurrences and nonlinearities (and unlike the original LHP), we enforce non-negative intensities by applying an affine projection followed by a rectifying transformation, similar to Mei and Eisner (2017). This is referred to as the “Proj. & Softplus” layer in Fig.˜1, and expressed as $\bm{\lambda}_{t}:=\mathbf{s}\odot\text{softplus}((\mathbf{W}\mathbf{u}_{t-}^{(L+1)}+\mathbf{b})\odot\mathbf{s}^{-1})$ for $t\geq 0$ and where $\odot$ is an element-wise product, $\mathbf{W}\in\mathbb{R}^{K\times H}$, and $\mathbf{b},\log(\mathbf{s})\in\mathbb{R}^{K}$. Then S2P2 can be trained by maximizing the log-likelihood of each sequence, Eq.˜2. Similar to other neural MTPPs, we use Monte Carlo estimation for the integral term $\int_{0}^{T}\lambda_{s}\mathrm{d}N_{s}$ (Mei and Eisner, 2017). Training the model requires computing intensities at both event times $t_{1:N}$ and at randomly sampled times $t\!\sim\!\mathcal{U}(0,T)$.
To compute intermediate intensities at these sampled points, we take advantage of the continuous-time nature of S2P2 and partially evolve the latent state through the system dynamics. To do so, we first compute the right limits of the hidden states at event times, $\mathbf{x}_{t_{i}}^{(1:L)}$, as described above (we can do this conditioning in logarithmic depth). Then, Eq.˜12 is applied, with no impulse, as no events are occurring at these intermediate points, to find $\mathbf{x}_{t}^{(1:L)}$. From there, the evolved hidden states are used to evaluate the model across depth to compute the intermediate intensity $\bm{\lambda}_{t}$. This operation is both efficient and can be done in constant complexity because intermediate evaluations are conditionally independent given the right limits at events. Crucially, there is no separate parametric decoding head, unlike, for instance, Mamba Hawkes processes (Gao et al., 2024) or transformer Hawkes processes (Zuo et al., 2020). Instead, this tying of the intensity to the model’s continuously evolving hidden states, more like the neural Hawkes processes (Mei and Eisner, 2017), makes S2P2 a continuous-time model and contributes to its enhanced expressivity (see Section˜5 and Section˜7 for more discussions). In Algorithms˜1, 2 and 3, we explicitly detail how to use a parallel scan to compute the sequence of right limits at events, how to then evolve those to compute left limits, and then how to subsequently compute the log-likelihood of the sequence.
4 Related Work